# Towards Thinking-Optimal Scaling of Test-Time Compute for LLM Reasoning

### v1

@@ 4 Thinking-Optimal Test-Time Scaling
Thus, we propose a Thinking-OPtimal Scaling (TOPS) strategy aiming to achieve more effective and efficient test-time scaling for LLM reasoning. We define a System-2 thinking-optimal response as the shortest correct response that the model can generate using System-2 thinking: fewer tokens may lead to wrong answer while more tokens causes overthinking. Then, our method includes three stages: Format Imitation, Reasoning Effort-Conditioned Generation, and Self-Improvement.
@@
Format Imitation First, we require a small set of o1-like responses for a cold start, enabling the model to learn the format of System-2 thinking patterns, including searching, reflecting, verification, backtracking, etc. Preliminary results in Section 3.2 and previous studies [22, 14] have shown that a small number of o1-like responses is sufficient for the model to effectively learn the format of System-2 reasoning. Such a small set of seed data can be manually human-written or generated by existing o1-like models. However, different from previous studies [22, 23] that use a fixed length distribution of seed samples (i.e., directly generated by existing o1-like models), which may not be a thinking-optimal distribution for our base model, we instead create the seed data containing responses under different reasoning efforts (i.e., different length distributions). Specifically, we define a small number of seed problems as $\mathcal{P}_{s}$, and our goal is to curate a seed dataset $\mathcal{D}_{s}=\mathcal{D}_{s_{1}}\cup\cdots\cup\mathcal{D}_{s_{n}}$, where $\mathcal{D}_{s_{i}}=\{(e_{i},x,y_{e_{i}})|x\sim\mathcal{P}_{s}\}$ represents the responses to seed problems under a specific reasoning effort $e_{i}$. This follows the same procedure as data curation for the tag model in Section 3.2. Then, we train the base model on this dataset to obtain the tag model that can apply different levels of reasoning effort to perform System-2 thinking on a given problem:
@@
$y_{e_{i}}\sim\pi(e_{i},x|\boldsymbol{\theta}_{tag}),\quad x\sim\mathcal{P}_{a},$ (2)
@@ 5.1 Experimental Settings
Base Model We perform our Thinking-Optimal Scaling strategy mainly on Qwen2.5-32B-Instruct, as it serves as an appropriate base model for exploration on test-time scaling according to previous works [27, 22, 23].
Datasets First, we directly use the model Qwen2.5-32B-Tag created in Section 3.2 as the tag model for reasoning effort-conditioned generation. As mentioned before, this tag model is trained on the seed data that contains 1.3K problems from a subset of NuminaMath, and totally 3.9K responses under three types of reasoning efforts generated by QwQ-32B-Preview. We then use this tag model to generate responses on an additional subset of NuminaMath containing extra 50K problems under different reasoning efforts. On each problem, we sample only 1 response for each reasoning effort, though we believe that performing multiple samplings could further enhance effectiveness. For each problem, we select the shortest correct response among all three responses to the problem as the thinking-optimal response. Finally, we incorporate the responses corresponding to low reasoning effort from the seed data into the above generated dataset, resulting in a thinking-optimal dataset of about 26K samples for self-improvement. We denote the self-improved model created by our method as Qwen2.5-32B-TOPS. We then evaluate the performance of our model on three typical math reasoning benchmarks: GSM8K, MATH500, and AIME2024, which are introduced in Section 3.1 and Section 3.2.
Training and Evaluation Details In the SFT stage, the learning rate is $1\times 10^{-5}$, the batch size is $96$, and the number of epochs is 2. In inference, the decoding temperature is $1.0$, the maximum generation length is 16,384. We report the average accuracy across 5 random seeds in each experiment. In the evaluation of MATH500, we also use gpt-4o to assist in identifying missed cases caused by format issues by following Gao et al. [8].
Baselines Besides the base model Qwen2.5-32B-Instruct, we compare our self-improved model with several existing o1-like models that are based on the same base model:
  • QwQ-32B-Preview: One of the most popular o1-like reasoning models developed by Qwen Team.
  • STILL-2-32B [22]: A System-2 thinking model trained on 3.9K challenging math examples generated by QwQ-32B-Preview and DeepSeek-R1-Lite [^4] https://api-docs.deepseek.com/news/news1120.
  • Sky-T1-32B-Preview [23]: The reasoning model trained on 17K examples generated by QwQ-32B-Preview, including 10K math examples.
  • Qwen2.5-32B-Random: After the reasoning effort-based generation by the tag model, we randomly select a correct solution on each problem rather than the shortest one to form a thinking-suboptimal dataset, and train the base model on this dataset.
@@ 5.2 Main Results
Qwen2.5-32B-Instruct ($\text{Temp}=0.0$) 95.91 295.01 84.20 0576.89 16.67 1407.43
Qwen2.5-32B-Instruct ($\text{Temp}=1.0$) 95.30 296.98 82.84 0555.65 14.67 0855.62
@@
The results of each model are displayed in Table 2. Besides the accuracy, we also report the number of CoT tokens used by each model on each dataset for reference. [^5] For STILL-2-32B and Sky-T1-32B-Preview, we only calculate the number of tokens in the thought part and do not include the tokens in the summary part.
@@
Regarding the reasoning efforts (i.e., the number of reasoning tokens) used by each model to solve different difficulty levels of tasks, we observe that Qwen2.5-32B-TOPS uses fewer tokens on easier tasks like GSM8K compared to other models, effectively mitigating the issue of overthinking [3]. On the other hand, it tends to spend more time thinking on harder problems such as AIME2024, demonstrating adaptive reasoning depths.
@@ 5.3 Results of Iterative Self-Improvement
To further enhance the reasoning performance of our model on challenging problems, we perform iterative self-improvement on Qwen2.5-32B-TOPS. Specifically, we select additional 4500 MATH problems [13] (which have not appeared in the previously used problems) and the problems from AIME1983-2023. On each problem, we sample 8 responses from Qwen2.5-32B-TOPS. Then, we select the shortest correct response among 8 responses as the chosen response. One iterative self-improvement approach is to further supervised fine-tune Qwen2.5-32B-TOPS on the dataset composed of all chosen responses (shortest correct responses), resulting in Qwen2.5-32B-TOPS-Iter-SFT. Besides, we can also perform preference optimization. Specifically, if there are responses with incorrect final answers, we select the longest incorrect response as the rejected response to improve reasoning capability. Additionally, we include preference pairs where the rejected response is the shortest wrong response if there exists a wrong response that is shorter than the shortest correct response, to avoid causing the model to underthink. After obtaining the preference dataset, we perform Direct Preference Optimization (DPO) [28] on Qwen2.5-32B-TOPS to get Qwen2.5-32B-TOPS-Iter-DPO. Detailed experimental settings are in Appendix C.
@@ 6 Conclusion
In this work, we explore a potential issue under current pursuit of test-time scaling. Through empirical analysis on mathematical reasoning tasks, we first demonstrate that overly long CoTs can negatively impact the model’s reasoning performance in certain domains, emphasizing the need for an optimal length distribution during CoT length scaling. To tackle this, we propose a Thinking-Optimal Scaling strategy, which leverages a small set of seed data to teach LLMs to adopt varying levels of reasoning effort to perform System-2 thinking. Then, our approach allows models to identify the shortest correct response for self-improvement, leading to a more efficient and effective System-2 thinking. Experimental results show that our self-improved and further iteratively self-improved models, based on Qwen2.5-32B-Instruct, outperform existing distillation-based o1-like models and achieve performance that is on par with QwQ-32B-Preview across various math benchmarks.
Impact Statement
@@
Appendix A Details on Estimating The Number of Tokens in Hidden CoT of o1-mini by Qwen2.5 Tokenizer
@@
Appendix B User Prompt for gpt-4o to Determine the Number of (Erroneous) Reasoning Rounds
@@
Appendix C Experimental Settings
C.1 Training Settings in Format Imitation
In the format imitation stage, we perform SFT on the base model Qwen2.5-32B-Instruct on a small subset of seed data containing 1.3K problems sampled from NuminaMath along with responses with varying lengths for each problem. The statistics of the seed data is shown in Table 1. In SFT stage, the learning rate is $1\times 10^{-5}$, the batch size is $32$, the number of epochs is $3$.
C.2 Training Settings in Self-Improvement
In the self-improvement stage, we perform SFT on Qwen2.5-32B-Instruct on the curated thinking-optimal dataset for 2 epochs. The learning rate is $1\times 10^{-5}$, and the batch size is 96.
C.3 Training Settings in Iterative Self-Improvement
In the iterative self-improvement stage, for Qwen2.5-32B-TOPS-Iter-SFT, the learning rate is $1\times 10^{-6}$, the batch size is 32, and we set the training epoch to 1. For Qwen2.5-32B-TOPS-Iter-DPO, the learning rate is $5\times 10^{-7}$, the batch size is 32, the training epoch is 3.
C.4 Evaluation Settings
@@
For all o1-like models, we set the decoding temperature to 1.0 and average the results over 5 random seeds for each evaluation experiment. The maximum generation length is 16,384.

### v2

@@ 4 Thinking-optimal test-time scaling
Thus, we propose a Thinking-OPtimal Scaling (TOPS) strategy aiming to achieve more effective and efficient test-time scaling for LLM reasoning. We define a System-2 thinking-optimal response as the shortest correct response that the model can generate using System-2 thinking: fewer tokens may lead to wrong answer while more tokens causes overthinking. Then, our method includes three stages: Format Imitation, Reasoning Effort-Conditioned Generation, and Self-Improvement. The illustration of our method is shown in Figure 6.
@@
Format Imitation First, we require a small set of o1-like responses for a cold start, enabling the model to learn the format of System-2 thinking patterns, including searching, reflecting, verification, backtracking, etc. Preliminary results in Section 3.2 and previous studies [25, 14] have shown that a small number of o1-like responses are sufficient for the model to effectively learn the format of System-2 reasoning. Such a small set of seed data can be manually human-written or generated by existing o1-like models. However, different from previous studies [25, 26] that use a fixed length distribution of seed samples (i.e., directly generated by existing o1-like models), which may not be a thinking-optimal distribution for our base model, we instead create the seed data containing responses under different reasoning efforts (i.e., different length distributions). Specifically, we define a small number of seed problems as $\mathcal{P}_{s}$, and our goal is to curate a seed dataset $\mathcal{D}_{s}=\mathcal{D}_{s_{1}}\cup\cdots\cup\mathcal{D}_{s_{n}}$, where $\mathcal{D}_{s_{i}}=\{(e_{i},x,y_{e_{i}})|x\sim\mathcal{P}_{s}\}$ represents the responses to seed problems under a specific reasoning effort $e_{i}$. This follows the same procedure as data curation for the tag model in Section 3.2. Then, we train the base model on this dataset to obtain the tag model that can apply different levels of reasoning effort to perform System-2 thinking on a given problem:
@@
$y_{e_{i}}\sim\pi(\cdot|e_{i},x;\boldsymbol{\theta}_{tag}),\quad x\sim\mathcal{P}_{a},$ (2)
@@ 5.1 Experimental settings
Base Model We mainly display and compare the results of performing Thinking-Optimal Scaling on Qwen2.5-32B-Instruct, as it serves as an appropriate base model for exploration on test-time scaling according to previous works [30, 25, 26]. To validate the generalizability of our approach, we also perform additional experiments on LLaMA3.1-8B-Instruct, and put the results in Section 5.4.
Datasets First, we directly use the model Qwen2.5-32B-Tag created in Section 3.2 as the tag model for reasoning effort-conditioned generation. As mentioned before, this tag model is trained on the seed data that contains 1.3K problems from a subset of NuminaMath, and totally 3.9K responses under three types of reasoning efforts generated by QwQ-32B-Preview. We then use this tag model to generate responses on an additional subset of NuminaMath containing extra 50K problems under different reasoning efforts. On each problem, we sample only 1 response for each reasoning effort, though we believe that performing multiple samplings could further enhance effectiveness. For each problem, we select the shortest correct response among all three responses to the problem as the thinking-optimal response. Finally, we incorporate the responses corresponding to low reasoning effort from the seed data into the above generated dataset, resulting in a thinking-optimal dataset of about 26K samples for self-improvement. We denote the self-improved model created by our method as Qwen2.5-32B-TOPS. We then evaluate the performance of our model on three typical math reasoning benchmarks: GSM8K, MATH500, and AIME2024. The additional results on general reasoning tasks can be found in Appendix J.
Training and Evaluation Details In the SFT stage, the learning rate is $1\times 10^{-5}$, the batch size is $96$, and the number of epochs is 2. In inference, the decoding temperature is $1.0$, the maximum generation length is 16,384. We report the average accuracy across 5 random seeds in each experiment. The standard deviation results are in Appendix K. In the evaluation of MATH500, we also use gpt-4o to assist in identifying missed cases caused by format issues [8].
Baselines Besides the base model Qwen2.5-32B-Instruct, we compare our self-improved model with several existing o1-like models that are based on the same base model: (1) QwQ-32B-Preview: One of the most popular o1-like reasoning models developed by Qwen Team. (2) STILL-2-32B [25]: A System-2 thinking model trained on 3.9K challenging math examples generated by QwQ-32B-Preview and DeepSeek-R1-Lite [^5] https://api-docs.deepseek.com/news/news1120. (3) Sky-T1-32B-Preview [26]: The reasoning model trained on 17K examples generated by QwQ-32B-Preview, including 10K math examples. (4) Qwen2.5-32B-Random: After the reasoning effort-based generation, we randomly select a correct solution on each problem rather than the shortest one to form a thinking-suboptimal dataset, and train the base model on this dataset.
@@ 5.2 Main results
Qwen2.5-32B-Instruct ($\text{Temp.}=0.0$) 95.91 295.01 84.20 0576.89 16.67 1407.43
Qwen2.5-32B-Instruct ($\text{Temp.}=1.0$) 95.30 296.98 82.84 0555.65 14.67 0855.62
@@
The results of each model are displayed in Table 3. Besides the accuracy, we also report the number of CoT tokens used by each model on each dataset. [^6] For STILL-2-32B and Sky-T1-32B-Preview, we only calculate the number of tokens in the thought part and do not include the tokens in the summary part.
@@
Regarding the reasoning efforts (i.e., the number of reasoning tokens) used by each model to solve different difficulty levels of tasks, we observe that Qwen2.5-32B-TOPS uses fewer tokens on easier tasks like GSM8K compared to other models, effectively mitigating the issue of overthinking [3]. On the other hand, it tends to spend more time thinking on harder problems such as AIME2024. The comparison of reasoning tokens used by different models across various domains reflects our model’s ability to exhibit adaptive reasoning depths.
@@ 5.3 Results of iterative self-improvement
To further enhance the reasoning performance of our model on challenging problems, we perform iterative self-improvement on Qwen2.5-32B-TOPS. Specifically, we select additional 4500 MATH problems [13] (which have not appeared in the previously used problems) and the problems from AIME1983-2023. On each problem, we sample 8 responses from Qwen2.5-32B-TOPS. Then, we select the shortest correct response among 8 responses as the chosen response. One iterative self-improvement approach is to further supervised fine-tune Qwen2.5-32B-TOPS on the dataset composed of all chosen responses (shortest correct responses), resulting in Qwen2.5-32B-TOPS-Iter-SFT. Besides, we can also perform preference optimization. Specifically, if there are responses with incorrect final answers, we select the longest incorrect response as the rejected response to improve reasoning capability. Additionally, we include preference pairs where the rejected response is the shortest wrong response if there exists a wrong response that is shorter than the shortest correct response, to avoid causing the model to underthink. After obtaining the preference dataset, we perform Direct Preference Optimization (DPO) [32] on Qwen2.5-32B-TOPS to get Qwen2.5-32B-TOPS-Iter-DPO. Detailed experimental settings are in Appendix E.
@@
5.4 Results on LLaMA3.1-8B-Instruct
Here, we display the results of performing our Thinking-Optimal Test-Time Scaling strategy on LLaMA3.1-8B-Instruct in Table 4. The experimental setups are consistent with that in the main experiments above. The results demonstrate the generalizability of our method on other model architectures.
Table 4: The self-improvement results on LLaMA3.1-8B-Instruct. In each setting, the underlined value represents the best result for System-1 thinking models, while the bold value indicates the best result for System-2 thinking models.
Model GSM8K MATH500 AIME2024
Accuracy #Tokens Accuracy #Tokens Accuracy #Tokens
   System-1 thinking models
LLaMA3.1-8B-Instruct ($\text{Temp.}=0.0$) 82.18 0262.23 47.00 1801.76 06.67 5506.30
LLaMA3.1-8B-Instruct ($\text{Temp.}=1.0$) 76.21 0233.08 39.60 0733.56 04.67 1691.88
   System-2 thinking models
LLaMA3.1-8B-Random-SFT 87.94 1051.05 60.52 3627.23 04.67 8165.69
LLaMA3.1-8B-TOPS-SFT 88.54 0571.10 61.28 3254.01 08.00 7392.59
@@ 6 Conclusion
In this work, we aim to explore a potential issue under current pursuit of test-time scaling. Through empirical analysis on mathematical reasoning tasks, we first demonstrate that overly long CoTs can negatively impact the model’s reasoning performance in certain domains, emphasizing the need for optimal CoT length scaling. To tackle this, we propose a Thinking-Optimal Scaling strategy, which first leverages a small set of seed data to teach LLMs to adopt varying levels of reasoning effort to perform System-2 thinking. Then, our approach allows models to identify the shortest correct response for self-improvement, leading to more efficient and effective System-2 thinking. Experimental results show that our self-improved and iteratively self-improved models, based on Qwen2.5-32B-Instruct, outperform existing distillation-based o1-like models and achieve comparable performance with QwQ-32B-Preview across various math benchmarks.
7 Limitations
There are some limitations in the current work: (1) Our analysis mainly focuses on the math reasoning domain, because math tasks offer relatively accurate and reliable performance verification. However, we believe investigating the potential impact of long CoTs in other domains is important, which can be a promising direction in future work. We also provide some preliminary results in the general reasoning domain in Appendix J. (2) This work primarily focuses on the SFT setting. While investigating effective SFT and distillation is very important and practical—as highlighted in concurrent studies [25, 38]—exploring the impact of CoT length in the reinforcement learning (RL) setting is another compelling direction. We believe that our findings can be extended to RL-based scaling. Specifically, since RL assigns positive rewards (e.g., 1.0) to all solutions with correct final answers, shorter correct solutions with fewer erroneous reasoning steps should be preferred over longer correct ones with more errors. Over-rewarding the latter may encourage the model to generate incorrect intermediate steps first and rely on its imperfect self-correction ability to correct erroneous steps, rather than producing the correct answer in one go.
Acknowledgements
We sincerely thank all the anonymous reviewers and (S)ACs for their valuable comments and thoughtful suggestions. This work was supported by The National Natural Science Foundation of China (No. 62376273 and No.U2436209) and Beijing Nova Program (No. 20240484568).
Appendix A Impact statement
@@
Appendix B Details on estimating the number of tokens in hidden CoT of o1-mini by Qwen2.5 tokenizer
@@
We also display the comparison results on the CoT token count returned by the o1-mini API and the number of tokens we estimated using the Qwen2.5 tokenizer in Table 5. As we can see, the numbers do not differ significantly. Thus, we use the estimation results from the Qwen2.5 tokenizer in the main text in order to make a fair comparison of the number of reasoning tokens used by different models.
Table 5: CoT token count returned from o1-mini API and estimated by Qwen2.5 tokenizer.
Method MATH500 AIME2024
o1-mini API 1122.07 4861.87
Qwen2.5 Tokenizer 1110.88 4972.08
Appendix C Prompts for generating reasoning responses under different reasoning efforts
We put the system prompts for QwQ-32B-Preview to generate responses under different reasoning efforts in Figure 7.
Appendix D User prompt for gpt-4o to determine the number of (erroneous) reasoning rounds
We display the user prompt for gpt-4o to help determine the number of total and erroneous reasoning rounds in Figure 8.
@@ Appendix D User prompt for gpt-4o to determine the number of (erroneous) reasoning rounds
Figure 8: User prompt for gpt-4o to determine the number of reasoning rounds.
Appendix E Experimental settings
E.1 Training settings for tag models
[…]
E.2 Training settings in format imitation
[…]
E.3 Training settings in self-improvement
[…]
E.4 Training settings in iterative self-improvement
[…]
E.5 Evaluation settings
@@
For all o1-like models, we set the decoding temperature to 1.0 and average the results over 5 random seeds for each evaluation experiment. The maximum generation length is 16,384. All evaluations are conducted on 4 $\times$ NVIDIA A100 80G.
Appendix F Breakdown results of tag models on MATH500
Table 6: Breakdown results of tag models on MATH500 categorized by problem difficulty levels.
Model Level-1 Level-2 Level-3 Level-4 Level-5
Acc. #Tokens Acc. #Tokens Acc. #Tokens Acc. #Tokens Acc. #Tokens
LLaMA3.1-8B-Tag-Low 92.09 1415.00 78.89 1702.68 68.38 2238.64 54.22 3386.63 32.24 4719.16
LLaMA3.1-8B-Tag-Medium 87.44 2123.17 79.11 2775.34 72.19 3541.26 57.81 4467.70 35.07 6267.16
LLaMA3.1-8B-Tag-High 87.44 2924.74 78.67 3561.78 71.62 4447.92 54.22 5242.71 32.69 6735.32
Qwen2.5-32B-Tag-Low 97.67 0673.82 95.33 0987.61 97.14 1098.93 90.31 1802.25 80.30 3330.60
Qwen2.5-32B-Tag-Medium 96.74 1757.15 96.22 1761.37 97.14 1983.42 90.94 2692.74 82.69 4344.62
Qwen2.5-32B-Tag-High 96.74 2077.88 95.33 2291.11 95.43 2479.56 89.84 3223.67 82.24 4659.30
We display the breakdown results of tag models on MATH500 under different difficulty levels in Table 6. The results also validate the claim that continuously increasing the reasoning effort can indeed bring adverse effects, especially on lower-level problems (e.g., 92.09 (Low) -> 87.44 (Medium/High) and 97.67 (Low) -> 96.74 (Medium/High) on Level 1 problems for LLaMA3.1-8B-Tag and Qwen2.5-32B-Tag separately).
Appendix G Response length distribution of QwQ-32B-Preview under different reasoning effort-based system prompts
Table 7: Raw length distributions (response length rankings) of different reasoning-effort prompted responses on QwQ-32B-Preview.
Effort Longest % Middle % Shortest %
Low 15.2 18.3 66.5
Medium 36.4 45.4 18.2
Hight 48.4 36.2 15.4
We put the initial response length distribution of different reasoning-effort prompted responses on QwQ-32B-Preview in Table 7. As we can see, in nearly 15% of cases, the low reasoning-effort prompted responses are even the longest among all three responses. This indicates that only using direct prompting is unreliable to explore the impact of CoT length on reasoning performance. Therefore, when training tag models, for each training problem, we reorder the responses to ensure that the response lengths under different reasoning efforts follow an consistently increasing order from low to medium to high. In such a controlled setting, we can fairly observe the impact of CoT length on reasoning performance.
Appendix H Performance of QwQ-32B-Preview and QwQ-32B when directly prompted with different reasoning effort-based prompts
Table 8: The performance of QwQ-32B-Preview and QwQ-32B on three benchmarks when directly prompting the model with various reasoning effort-based prompts.
Reasoning Effort GSM8K MATH500 AIME2024
Accuracy #Tokens Accuracy #Tokens Accuracy #Tokens
   QwQ-32B-Preview
Low 93.95 00418.26 90.24 01592.29 42.00 05152.97
Medium 93.56 00844.32 89.16 02356.29 44.00 06413.97
High 92.78 01112.97 88.36 02684.87 41.33 06678.50
   QwQ-32B
Low 96.56 00568.49 95.76 02532.82 74.00 11390.70
Medium 96.50 01152.86 96.08 03708.79 78.00 12969.53
High 96.36 01752.29 95.80 04198.67 80.67 13194.29
Here, we display the performance of QwQ-32B-Preview [30] and QwQ-32B [31] on all three evaluation benchmarks when directly prompted with different reasoning effort-based prompts. Since QwQ-32B includes both reasoning and summary components in its responses and typically generates much longer CoTs, we extended its max_seq_len to 32K to allow for sufficient reasoning and summarization. The full results are in Table 8. These results re-validate our main claim: longer CoTs may not necessarily lead to better performance.
Appendix I Results of fine-tuning on samples after filtering out solutions with erroneous steps
We conduct additional experiments on supporting the claim that including more erroneous rounds in the training stage can lead to greater adverse effects. Specifically, we filter out those solutions that contain erroneous steps (identified by GPT-4.1) under the high reasoning effort, and fine-tune Qwen2.5-32B-Instruct on the filtered data, which yields Qwen2.5-32B-Tag-High-Filtered. The results are in Table 9.
Table 9: The performance of the model trained on data with high reasoning effort after filtering out all solutions that are identified to contain erroneous steps.
Model GSM8K MATH500 AIME2024
Accuracy #Tokens Accuracy #Tokens Accuracy #Tokens
Qwen2.5-32B-Tag-High 93.31 1820.17 90.56 3185.75 41.33 8753.87
Qwen2.5-32B-Tag-High-Filtered 94.87 1478.10 90.00 2783.98 36.33 8049.29
We have some interesting findings: (1) After removing samples containing erroneous rounds, the trained model produces much shorter CoTs. The reason is that the solutions removed are usually very lengthy as they contain more self-reflections and self-corrections, thus the average length of the dataset after removing these samples is much shorter. (2) Removing samples containing erroneous rounds brings significant performance improvement on GSM8K, while causing certain performance degradation on harder benchmarks MATH500 and AIME2024. The reason is that after removing those samples, the model cannot learn to effectively perform deeper thinking such as correcting the errors it could make in previous rounds. GSM8K is relatively simple, so the model does not need to engage in excessive reflection and correction. Therefore, removing these behaviors actually improves model performance. On more challenging datasets, the advantages of longer reasoning chains that include reflections and corrections become apparent. Moreover, by comparing with the results in Figure 5 in the main text, we find that instead of directly removing correct solutions that contain erroneous steps, a more effective approach is to apply loss masking specifically to the wrong steps. The latter strategy allows the model to retain the ability to learn how to correct previous mistakes, without explicitly learning from the erroneous steps themselves.
Appendix J Results on general reasoning tasks
Here, we additionally prompt QwQ-32B-Preview to generate responses under different reasoning effort-based prompts on a subset of the WebInstruct-verified dataset [23], and curate a seed general reasoning dataset that contains correct responses with varying reasoning efforts. We then fine-tune Qwen2.5-7B-Instruct [29] on the seed dataset to get Qwen2.5-7B-Tag-General, and evaluate its performance on MMLU-Pro [41] and GPQA-Diamond [33] using the same reasoning effort-conditioned prompting strategy as in Section 3.2. We only sample once for each prompt considering the evaluation cost. The results are displayed in Table 10. The findings are consistent with the results in the math reasoning domain that excessive scaling with longer CoTs can bring negative effects to the model’s performance in general reasoning tasks. Also, the model needs more reasoning effort to perform well on the more challenging task GPQA-Diamond.
Table 10: The performance of Qwen2.5-7B-based models on MMLU-Pro and GPQA-Diamond
Model MMLU-Pro GPQA-Diamond
Accuracy #Tokens Accuracy #Tokens
   System-1 thinking models
Qwen2.5-7B-Instruct ($\text{Temp.}=0.0$) 52.46 0401.38 34.85 0592.73
Qwen2.5-7B-Instruct ($\text{Temp.}=1.0$) 51.49 0379.84 33.84 0537.41
   Tag models
Qwen2.5-7B-Tag-General-Low 56.00 1674.60 31.82 2808.88
Qwen2.5-7B-Tag-General-Medium 55.92 2341.27 36.87 3931.13
Qwen2.5-7B-Tag-General-High 55.81 2632.05 32.83 4238.11
   System-2 thinking models
Qwen2.5-7B-Random-General 55.86 1960.87 34.34 3385.70
Qwen2.5-7B-TOPS-General (ours) 56.50 1788.74 38.38 3083.76
Following the above setup, we then perform our thinking-optimal scaling on a held-out set of the WebInstruct-verified dataset to get Qwen2.5-7B-TOPS-General and perform random scaling to get Qwen2.5-7B-Random-General. The evaluation results in Table 10 show that our method can also work well for general reasoning tasks.
Appendix K Standard deviation results
Table 11: The detailed standard deviation results on LLaMA3.1-8B-Instruct.
Model GSM8K MATH500 AIME2024
   System-1 thinking models
LLaMA3.1-8B-Instruct ($\text{Temp.}=0.0$) 82.18 ($\pm$ 0.00) 47.00 ($\pm$ 0.00) 06.67 ($\pm$ 0.00)
LLaMA3.1-8B-Instruct ($\text{Temp.}=1.0$) 76.21 ($\pm$ 0.66) 39.60 ($\pm$ 0.84) 04.67 ($\pm$ 2.98)
   System-2 thinking models
LLaMA3.1-8B-Random-SFT 87.94 ($\pm$ 0.33) 60.52 ($\pm$ 1.38) 04.67 ($\pm$ 1.83)
LLaMA3.1-8B-TOPS-SFT 88.54 ($\pm$ 0.26) 61.28 ($\pm$ 0.73) 08.00 ($\pm$ 1.83)
Table 12: The detailed standard deviation results on Qwen2.5-32B-Instruct.
Model GSM8K MATH500 AIME2024
System-1 thinking models
Qwen2.5-32B-Instruct ($\text{Temp.}=0.0$) 95.91 ($\pm$ 0.00) 84.20 ($\pm$ 0.00) 16.67 ($\pm$ 0.00)
Qwen2.5-32B-Instruct ($\text{Temp.}=1.0$) 95.30 ($\pm$ 0.36) 82.84 ($\pm$ 0.65) 14.67 ($\pm$ 5.06)
System-2 thinking models
QwQ-32B-Preview 95.23 ($\pm$ 0.39) 92.02 ($\pm$ 0.87) 45.33 ($\pm$ 3.80)
STILL-2-32B 95.47 ($\pm$ 0.26) 91.40 ($\pm$ 0.80) 45.33 ($\pm$ 6.06)
Sky-T1-32B-Preview 94.82 ($\pm$ 0.33) 89.48 ($\pm$ 0.78) 35.33 ($\pm$ 2.98)
Qwen2.5-32B-Random 95.00 ($\pm$ 0.31) 90.16 ($\pm$ 0.95) 39.33 ($\pm$ 2.79)
Qwen2.5-32B-TOPS (ours) 95.82 ($\pm$ 0.15) 91.48 ($\pm$ 0.59) 43.33 ($\pm$ 2.36)
Qwen2.5-32B-TOPS-Iter-SFT (ours) 95.45 ($\pm$ 0.14) 90.76 ($\pm$ 0.71) 44.00 ($\pm$ 3.65)
Qwen2.5-32B-TOPS-Iter-DPO (ours) 95.80 ($\pm$ 0.13) 91.60 ($\pm$ 0.62) 46.00 ($\pm$ 3.65)
We put the detailed standard deviation results in Table 11 and Table 12 for reference.